OpenStudy (anonymous):

What is the y-intercept of the graph of 3x – 5y = 15 ? (0, –5) (0, 5) (0, –3) (0, 3)

OpenStudy (anonymous):

third one

OpenStudy (anonymous):

to find the y intercept, set x to be zero and solve for y: 3(0)-5y=15 -5y=15 y=-3, thus we have the y intercerpt as being: (0,-3)